Minutes


Talk time : 11:50, Duration : 00:20

 
1) Run14 production status:
 - Hao's tracking  changes committed by Dmitri
 - still need: a) IST fastsim/hit maker in the library and b) finish TPC calibrations [Mustafa/Alex's work brought this down to ~250um level. To be discussed in todays TPC meeting]. We may also want to [Xin] monitor the TPC space-charge corrections with the HFT. 

2) Run15 preps
 - Next Tues 1/13 STAR shifts start and we should be able to start taking B=0 cosmics
 - QA
      *  Shushu prepares the offline QA machinery
      * This includes the PXL masking. For this year the Warsaw group [Leszek] with Mike will do the daily review/monitoring
      * IST group is discussing/organizing this for the subdetector
      * SSD in QA - Jonathan to update plots/verify they are included in chain
      * Daniel to contact Gene to make sure the QA plots for Shift-crew are included in this run
 - Survey data of new PXL in Db...in progress. For IST last year's numbers are assumed to be fine.
 - We should run the Alignment chain the moment the cosmic data is taken and processed [priority item] [Alex, Long and Mike]